Integrals Giving Inverse Trigonometric Functions


∫(xdx)/√(4-x²)


1
Expert's answer
2021-04-15T07:21:40-0400

"\int{\frac{x}{\sqrt{4-x^2}}dx}"

"\ let\ u=\ \sqrt{4-x^2}"

"\ du=\ -\frac{x}{\sqrt{4-x^2}}dx"

"\ \int{-du\ }"

"-u + c"

"-\ \sqrt{4-x^2}+c"